OpenStudy (anonymous):

the average amount of savings of 10 students is $60. if 3 of the students have no savings at all, and each of the others has at least $25, including john, who has exactly $130, what is the largest amount in dollars that any one student can have?

OpenStudy (ganpat):

number of students = 10. Avg savings = 60$ So total money = 60 * 10 = 600$ John has 130$ 3 students = 0$ so 3 + John + 6students = 600 0 + 130 + 6students = 600 6 students = 470.. this all 6 students has min 25.. so 6 * 25 = 150.. 470 - 150 = 320.. so maximum one person can have is 320 + 25 = 345$.. i guess, do u have any options ??

OpenStudy (amistre64):

i was thinking, and it could be wrong, to try to undo the process from a known point n/10 = 60 n=60*10 = 600 n = 0+0+0+130 + >25 +>25 + >25 600 = 130 + something 600 - 130 = 470 470 = >25 + >25 + >25 = 25+a + 25 + b + 25+ c 470 - 75 = a+b+c 395 = a+b+c

OpenStudy (amistre64):

if we assume that a b or c has to have at least 1 dollar more than 25 that gets me: 395 = 1+1+largest amount possible 393

OpenStudy (amistre64):

0+0+0+130+26+26+393 = 575 .... got lost somewhere in the mathing lol i ended up with 7 students :/

OpenStudy (ganpat):

@amistre64 : i doubt ur method..

OpenStudy (amistre64):

0 + 0 + 0 + 130 + 26 + 26 + 26 + 26 + 26 + L = 600 there, thats 10 peoples; with 3 at 0, john at 130, and everyone else having more than 25 dollars :) oh, i doubt it too

OpenStudy (amistre64):

340 is what i end up with ....

OpenStudy (ganpat):

0 + 0 + 0 + 130 + 25 + 25 + 25 + 25 + 25 + x = 600 x = 345..

OpenStudy (ganpat):

y u took 26 ?? they say atleast 25.. so it can 25 as well ??

OpenStudy (amistre64):

at least $25 .... oh, i read it wrong :)

OpenStudy (amistre64):

then yes, 345 is good lol

OpenStudy (amistre64):

at any rate, you no longer ahve to guess if that was right ....

OpenStudy (ganpat):

h aha , just wanted to recheck.. if it was wrong, penguinn would have been in trouble.. :D

OpenStudy (ganpat):

@penguinn : so did u understood ??

OpenStudy (anonymous):

yea ty. :)

OpenStudy (ganpat):

great !!